Math 150 Find all of the real zeros and write the polynomial as a product of linear factors. Show all work.
x,cubed - 11x,squared - 15x + 325

Answer by stanbon(75887)

Math 150 Find all of the real zeros and write the polynomial as a product of linear factors. Show all work.
x,cubed - 11x,squared - 15x + 325
------------
x^3 - 11x^2 - 15x + 325
----
I graphed and found a zero at x = -5
---
Use synthetic division to find the other zeros:
-5)....1....-11....-15...325
........1.....-16....65..|..0
----
Quotient: x^2-16x+65
x = [16 +- sqrt(16^2-4*65)]/2
----
x = [16 +- sqrt(-4)]/2
----
x = 8+(1/2)2i or x = 8-(1/2)2i
Note: These are not Real Number zeroes
----
f(x) = (x+5)(x^2-16x+65)
===========================
Cheers,
Stan H.
=============